Circular No. 8759

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

RINGS OF SATURN (R/2006 S 1, R/2006 S 2, R/2006 S 3, R/2006 S 4)

C. C. Porco, CICLOPS, Space Science Institute, Boulder; and the Cassini Imaging Science Team report the sighting of several new rings around Saturn. One new diffuse ring, R/2006 S 1, has been found at an orbital distance of ∼ 151500 km, coincident with the co-orbital satellites, Saturn X (Janus) and XI (Epimetheus) with a radial width of ∼ 5000 km. Another new diffuse ring, R/2006 S 2, has been found coincident with the tiny moon Saturn XXXIII (Pallene; IAUC 8389), at an orbital distance of ∼ 212000 km and radial width of ∼ 2500 km. Two other newly sighted rings have been seen within the Cassini division. R/2006 S 3 is located in the outer gap of the Cassini division at an orbital distance of 119930 km, with a radial width of ∼ 50 km. Finally, a very narrow, discontinuous, and tenuous ringlet, R/2006 S 4, has been seen in between two broad bands in the Cassini division, at an orbital distance of 118960 km; its width is ∼ 6 km. R/2006 S 3 appears to have substantial numbers of small particles as it brightens considerably at high-phase-viewing geometries.

SUPERNOVAE 2006iv, 2006iw, 2006ix, 2006iy, 2006iz, 2006ja An apparent supernova (2006iv) has been discovered on unfiltered CCD

images by Grzegorz Duszanowicz (Akersberga, Sweden, 0.32-m f/3.1 reflector); he measured the position of the new object to be α = 11^{h}48^{m}12^{s}.35, δ = +54^{o}59′14′′.6 (equinox 2000.0), which is 7′′ east and 18′′ south of the center of UGC 6774. Approximate magnitudes for 2006iv from Duszanowicz’s unfiltered images: Jan. 3.145 UT, [18.5; Sept. 13.989, [18.0; 22.92, 16.7; Oct. 5.875, 15.9 (discovery); 7.87, 15.9.

Five additional faint supernovae (2006iw–2006ja; all in the magnitude range g = 20.4–22.6) have been reported by J. Frieman on behalf of the SDSS II collaboration (details on CBETs 663 and 667); 2006iw and 2006ix are type-II supernovae, 2006iy and 2006iz are type-Ia supernovae, and 2006ja is a probable type-Ia supernova.

COMET C/2006 P1 (McNAUGHT) Improved orbital elements (cf. IAUC 8737) from MPC 57794:

T = 2007 Jan. 12.8217 TT  ω = 155.^{o}9746
Ω = 267.4139 }2000.0
q = 0.170773 AU  i = 77.8512
2006 October 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green